### v1

@@
The dependence of (980) production on the final-state charged-particle multiplicity is reported for proton–proton (pp) collisions at the centre-of-mass energy, 13 TeV. The production of (980) is measured with the ALICE detector via the decay channel in a midrapidity region of 0.5. The evolution of the integrated yields and mean transverse momentum of f0(980) as a function of charged-particle multiplicity measured in pp at 13 TeV follows the trends observed in pp at 5.02 TeV and in proton–lead (p–Pb) collisions at 5.02 TeV. Particle yield ratios of (980) to and (892)0 are found to decrease with increasing charged-particle multiplicity. These particle ratios are compared with calculations from the canonical statistical thermal model as a function of charged-particle multiplicity. The thermal model calculations provide a better description of the decreasing trend of particle ratios when no strange or antistrange quark composition for f0(980) is assumed, which suggests that the tetraquark interpretation of the f0(980) is disfavored.
@@ 1 Introduction
Light scalar mesons, one of the most frequently revisited hadron classes, consist of light-flavoured quarks and are characterized by quantum numbers and masses below 2 GeV/. The study of light scalar mesons, including their internal structure, properties such as decay width, and formation mechanisms, is an actively explored research area [9, 10, 11]. The is a light scalar meson and has a large mass ( MeV/) and a relatively small particle width ( 40–70 MeV/) [12, 13]. Given these characteristics, the can currently be experimentally identified compared to other light scalar states, such as and , where the threshold mass is close to the invariant mass. Theoretical interpretations of the elusive internal structure of this particle suggest that can either be a conventional meson () [14], a compact tetraquark [15], or a molecule [16]. These competing scenarios remain the subject of intense theoretical debate [12].
@@
QCD predicts the formation of a deconfined and strongly interacting matter, called the quark–gluon plasma (QGP), under conditions of high temperature and high energy density [17, 18, 19, 20]. As the system cools down, the near-perfect fluidity of QGP collapses, confining quarks and gluons into hadrons upon crossing its critical temperature. As the system cools below the critical temperature, all inelastic interactions stop, confining partons into hadrons, where the relative particle composition is expected to be fixed at the chemical freeze-out stage. In heavy-ion collisions, the thermally produced hadron gas can be modelled according to a grand canonical ensemble of particle species, considering the conservation of baryon number, electric charge, strangeness, and other quantum numbers on average. Statistical hadronisation modelling of particle production can be extended to pp and p–Pb collisions by employing a canonical approach. An implementation of the statistical model is the canonical statistical model with undersaturation of strangeness, CSM [21]. The CSM accounts for the system-size-dependent hadrochemistry at vanishing baryon density with local conservation of electric charge, baryon density, as well as strangeness. It can, therefore, predict the relative particle composition while suppressing the production of strangeness in small collision systems. Although CSM does not account for the hadronic phase or the rescattering of decay products from short-lived resonances, it remains a powerful tool for investigating strangeness production. In particular, such an approach can provide insight into whether certain hadrons include strange quarks. Due to the exotic nature of and its production mechanisms, it has not yet been implemented in most QCD-inspired general-purpose Monte Carlo models utilised in the field. Therefore, the CSM offers a suitable framework for studying the strangeness-related properties of the . For instance, the enhancement of ratio with increasing multiplicity can indicate the presence of strange quarks inside the . On the other hand, the absence of late hadronic interactions in the statistical hadronisation model results in the overestimation of resonance yields, especially in large collision systems. Therefore, the model overestimates the observed (980)/ ratio. One open point, which is relevant to a precise data-to-model comparison [22], is the knowledge of the (980) branching ratios (B.R.). Only few measurements of the B.R. of the decay channel [23, 24, 25] are available. The CELO [23] and BaBar [24] measurements of the ratio of B.R. are compatible with zero within 3 units of their uncertainties, where the zero value of the ratio indicates no decay. Such a measurement of the B.R. is relevant because it directly affects the corrected yields. However, after chemical freeze-out, the system still evolves through a hadronic phase, during which elastic interactions and resonance dynamics play an important role.
@@ 5 Results
Figure 4 shows the particle yield ratio of to charged pions as a function of charged-particle multiplicity in different collision systems. The decreases with increasing multiplicity across different collision systems. The experimental data are compared with theoretical calculations from the CSM [21]. The CSM calculation with the assumption of 2 for expects to increase with increasing multiplicity because it assumes a specific quark content for the , including strange quarks. On the other hand, the CSM calculation with the assumption of 0 for qualitatively describes the decreasing trend, which is observed in experimental data. However, a quantitative description is not achieved as the model is consistent only with the point in the highest-multiplicity events, and the model provides an underestimation of 39% of the particle yield ratio in the lowest-multiplicity events. The model provides good description of particle yield ratio in low multiplicity events and overestimates the ratio in high multiplicity events for other hadron species [21]. In both the and scenarios, an additional decrease with increasing multiplicity can be expected if the rescattering effect is considered, and in the scenario, the increase could be converted to a decrease. The model prediction for the scenario is closer to the data.
Figure 5 shows the particle yield ratio of to K∗(892)0 as a function of charged-particle multiplicity in different collision systems. The K∗(892)0 decreases with increasing multiplicity in different collision systems. The CSM calculation with the assumption of 0 also predicts the decreasing trend of the particle yield ratio, whereas the calculation with the the scenario does not. However, with the scenario, the model underestimates the particle yield ratio by about 35% throughout the multiplicity range, which is of a similar magnitude to the underestimation observed in the / ratio. Because rescattering affects both the and (892)0 resonances, its effects are expected to partially cancel out in the /(892)0 yield ratio. However, despite their similar lifetimes, they differ in their decay products, where and (892)0 decay into and K, respectively. The rescattering effect for can be stronger than for (892)0 when the scattering cross section is greater than . Similar to ratio, the model prediction for (892)0 with the scenario is closer to the data.
@@ Acknowledgements
The ALICE Collaboration would like to thank all its engineers and technicians for their invaluable contributions to the construction of the experiment and the CERN accelerator teams for the outstanding performance of the LHC complex. The ALICE Collaboration gratefully acknowledges the resources and support provided by all Grid centres and the Worldwide LHC Computing Grid (WLCG) collaboration. The ALICE Collaboration acknowledges the following funding agencies for their support in building and running the ALICE detector: A. I. Alikhanyan National Science Laboratory (Yerevan Physics Institute) Foundation (ANSL), State Committee of Science and World Federation of Scientists (WFS), Armenia; Austrian Academy of Sciences, Austrian Science Fund (FWF): [M 2467-N36] and Nationalstiftung für Forschung, Technologie und Entwicklung, Austria; Ministry of Communications and High Technologies, National Nuclear Research Center, Azerbaijan; Rede Nacional de Física de Altas Energias (Renafae), Financiadora de Estudos e Projetos (Finep), Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) and The Sao Paulo Research Foundation (FAPESP), Brazil; Bulgarian Ministry of Education and Science, within the National Roadmap for Research Infrastructures 2020-2027 (object CERN), Bulgaria; Ministry of Education of China (MOEC) , Ministry of Science & Technology of China (MSTC) and National Natural Science Foundation of China (NSFC), China; Ministry of Science and Education and Croatian Science Foundation, Croatia; Centro de Aplicaciones Tecnológicas y Desarrollo Nuclear (CEADEN), Cubaenergía, Cuba; Ministry of Education, Youth and Sports of the Czech Republic, Czech Republic; The Danish Council for Independent Research | Natural Sciences, the VILLUM FONDEN and Danish National Research Foundation (DNRF), Denmark; Helsinki Institute of Physics (HIP), Finland; Commissariat à l’Energie Atomique (CEA) and Institut National de Physique Nucléaire et de Physique des Particules (IN2P3) and Centre National de la Recherche Scientifique (CNRS), France; Bundesministerium für Bildung und Forschung (BMBF) and GSI Helmholtzzentrum für Schwerionenforschung GmbH, Germany; General Secretariat for Research and Technology, Ministry of Education, Research and Religions, Greece; National Research, Development and Innovation Office, Hungary; Department of Atomic Energy Government of India (DAE), Department of Science and Technology, Government of India (DST), University Grants Commission, Government of India (UGC) and Council of Scientific and Industrial Research (CSIR), India; National Research and Innovation Agency - BRIN, Indonesia; Istituto Nazionale di Fisica Nucleare (INFN), Italy; Japanese Ministry of Education, Culture, Sports, Science and Technology (MEXT) and Japan Society for the Promotion of Science (JSPS) KAKENHI, Japan; Consejo Nacional de Ciencia (CONACYT) y Tecnología, through Fondo de Cooperación Internacional en Ciencia y Tecnología (FONCICYT) and Dirección General de Asuntos del Personal Academico (DGAPA), Mexico; Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO), Netherlands; The Research Council of Norway, Norway; Pontificia Universidad Católica del Perú, Peru; Ministry of Science and Higher Education, National Science Centre and WUT ID-UB, Poland; Korea Institute of Science and Technology Information and National Research Foundation of Korea (NRF), Republic of Korea; Ministry of Education and Scientific Research, Institute of Atomic Physics, Ministry of Research and Innovation and Institute of Atomic Physics and Universitatea Nationala de Stiinta si Tehnologie Politehnica Bucuresti, Romania; Ministerstvo skolstva, vyskumu, vyvoja a mladeze SR, Slovakia; National Research Foundation of South Africa, South Africa; Swedish Research Council (VR) and Knut & Alice Wallenberg Foundation (KAW), Sweden; European Organization for Nuclear Research, Switzerland; Suranaree University of Technology (SUT), National Science and Technology Development Agency (NSTDA) and National Science, Research and Innovation Fund (NSRF via PMU-B B05F650021), Thailand; Turkish Energy, Nuclear and Mineral Research Agency (TENMAK), Turkey; National Academy of Sciences of Ukraine, Ukraine; Science and Technology Facilities Council (STFC), United Kingdom; National Science Foundation of the United States of America (NSF) and United States Department of Energy, Office of Nuclear Physics (DOE NP), United States of America. In addition, individual groups or members have received support from: Czech Science Foundation (grant no. 23-07499S), Czech Republic; FORTE project, reg. no. CZ.02.01.01/00/22_008/0004632, Czech Republic, co-funded by the European Union, Czech Republic; European Research Council (grant no. 950692), European Union; Deutsche Forschungs Gemeinschaft (DFG, German Research Foundation) “Neutrinos and Dark Matter in Astro- and Particle Physics” (grant no. SFB 1258), Germany; ICSC - National Research Center for High Performance Computing, Big Data and Quantum Computing and FAIR - Future Artificial Intelligence Research, funded by the NextGenerationEU program (Italy).

### v2

@@
The dependence of (980) production on the final-state charged-particle multiplicity is reported for proton–proton (pp) collisions at the centre-of-mass energy, 13 TeV. The production of (980) is measured with the ALICE detector via the decay channel in a midrapidity region of 0.5. The evolution of the integrated yields and mean transverse momentum of f0(980) as a function of charged-particle multiplicity measured in pp at 13 TeV follows the trends observed in pp at 5.02 TeV and in proton–lead (p–Pb) collisions at 5.02 TeV. Particle yield ratios of (980) to and (892)0 are found to decrease with increasing charged-particle multiplicity. These particle ratios are compared with calculations from the canonical statistical thermal model as a function of charged-particle multiplicity. The thermal model calculations provide a better description of the decreasing trend of particle ratios when no strange or antistrange quark composition for f0(980) is assumed, which suggests that the data do not support significant hidden strangeness in the .
@@ 1 Introduction
Light scalar mesons, one of the most frequently revisited hadron classes, consist of light-flavoured quarks and are characterized by quantum numbers and masses below 2 GeV/. The study of light scalar mesons, including their internal structure, properties such as decay width, and formation mechanisms, is an actively explored research area [9, 10, 11]. The is a light scalar meson and has a large mass ( MeV/) and a relatively small particle width ( 40–70 MeV/) [12, 13]. Given these characteristics, the can currently be experimentally identified compared to other light scalar states, such as and , where the threshold mass lies close to the pole mass. Theoretical interpretations of the elusive internal structure of this particle suggest that can either be a conventional meson () [14], a compact tetraquark [15], or a molecule [16]. These competing scenarios remain the subject of intense theoretical debate [12].
@@
QCD predicts the formation of a deconfined and strongly interacting matter, called the quark–gluon plasma (QGP), under conditions of high temperature and high energy density [17, 18, 19, 20]. As the system cools down, the near-perfect fluidity of QGP collapses, confining quarks and gluons into hadrons upon crossing its critical temperature. As the system cools below the critical temperature, all inelastic interactions stop, confining partons into hadrons, where the relative particle composition is expected to be fixed at the chemical freeze-out stage. In heavy-ion collisions, the thermally produced hadron gas can be modelled according to a grand canonical ensemble of particle species, considering the conservation of baryon number, electric charge, strangeness, and other quantum numbers on average. Statistical hadronisation modelling of particle production can be extended to pp and p–Pb collisions by employing a canonical approach. An implementation of the statistical model is the canonical statistical model with undersaturation of strangeness, CSM [21]. The CSM accounts for the system-size-dependent hadrochemistry at vanishing baryon density with local conservation of electric charge, baryon density, as well as strangeness. It can, therefore, predict the relative particle composition while suppressing the production of strangeness in small collision systems. Although CSM does not account for the hadronic phase or the rescattering of decay products from short-lived resonances, it remains a powerful tool for investigating strangeness production. In particular, such an approach can provide insight into whether certain hadrons include strange quarks. Due to ongoing studies on the possible exotic nature of and its production mechanisms, it has not yet been implemented in most QCD-inspired general-purpose Monte Carlo models utilised in the field. Therefore, the CSM offers a suitable framework for studying the strangeness-related properties of the .
In this context, two limiting hypotheses for the strangeness content of the are considered within the CSM. The first case, , corresponds to a state without valence strange quarks, consistent with light-quark configurations such as mixtures of and components. The second case, , represents scenarios in which the carries hidden strangeness through the presence of an pair. This assignment covers different possible internal structures, including -rich quark–antiquark states, four-quark configurations, or molecular components. Here, does not specify a particular internal arrangement, but simply encodes the presence of hidden strangeness that determines the model prediction. For instance, the enhancement of ratio with increasing multiplicity can indicate the presence of strange quarks inside the . On the other hand, the absence of late hadronic interactions in the statistical hadronisation model results in the overestimation of resonance yields, especially in large collision systems. Therefore, the model overestimates the observed (980)/ ratio. One open point, which is relevant to a precise data-to-model comparison [22], is the knowledge of the (980) branching ratios (B.R.). Only few measurements of the B.R. of the decay channel [23, 24, 25] are available. The CLEO [23] and BaBar [24] measurements of the ratio of B.R. are compatible with zero within 3 units of their uncertainties, where a ratio consistent with zero indicates the absence of the decay channel. Such a measurement of the B.R. is relevant because it directly affects the corrected yields. However, after chemical freeze-out, the system still evolves through a hadronic phase, during which elastic interactions and resonance dynamics play an important role.
@@ 5 Results
Figure 4 shows the particle yield ratio of to charged pions as a function of charged-particle multiplicity in different collision systems. The decreases with increasing multiplicity across different collision systems. The experimental data are compared with theoretical calculations from the CSM [21]. The CSM calculation with the assumption of 2 for expects to increase with increasing multiplicity because it assumes a specific quark content for the , including strange quarks. On the other hand, the CSM calculation with the assumption of 0 for qualitatively describes the decreasing trend, which is observed in experimental data. However, a quantitative description is not achieved as the model is consistent only with the point in the highest-multiplicity events, and the model provides an underestimation of 39% of the particle yield ratio in the lowest-multiplicity events. The model provides good description of particle yield ratio in low-multiplicity events and overestimates the ratio in high-multiplicity events for other hadron species [21]. In both the and scenarios, an additional decrease with increasing multiplicity can be expected if the rescattering effect is considered, and in the scenario, the increase could be converted to a decrease. The model prediction for the scenario is closer to the data.
Figure 5 shows the particle yield ratio of to K∗(892)0 as a function of charged-particle multiplicity in different collision systems. The K∗(892)0 decreases with increasing multiplicity in different collision systems. The CSM calculation with the assumption of 0 also predicts the decreasing trend of the particle yield ratio, whereas the calculation with the scenario does not. However, with the scenario, the model underestimates the particle yield ratio by about 35% throughout the multiplicity range, which is of a similar magnitude to the underestimation observed in the / ratio. Because rescattering affects both the and (892)0 resonances, its effects are expected to partially cancel out in the /(892)0 yield ratio. However, despite their similar lifetimes, they differ in their decay products, where and (892)0 decay into and K, respectively. The rescattering effect for can be stronger than for (892)0 when the scattering cross section is greater than . Similar to ratio, the model prediction for (892)0 with the scenario is closer to the data.
@@ Acknowledgements
The ALICE Collaboration would like to thank all its engineers and technicians for their invaluable contributions to the construction of the experiment and the CERN accelerator teams for the outstanding performance of the LHC complex. The ALICE Collaboration gratefully acknowledges the resources and support provided by all Grid centres and the Worldwide LHC Computing Grid (WLCG) collaboration. The ALICE Collaboration acknowledges the following funding agencies for their support in building and running the ALICE detector: A. I. Alikhanyan National Science Laboratory (Yerevan Physics Institute) Foundation (ANSL), State Committee of Science and World Federation of Scientists (WFS), Armenia; Austrian Academy of Sciences, Austrian Science Fund (FWF): [M 2467-N36] and Nationalstiftung für Forschung, Technologie und Entwicklung, Austria; Ministry of Communications and High Technologies, National Nuclear Research Center, Azerbaijan; Rede Nacional de Física de Altas Energias (Renafae), Financiadora de Estudos e Projetos (Finep), Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP) and The Sao Paulo Research Foundation (FAPESP), Brazil; Bulgarian Ministry of Education and Science, within the National Roadmap for Research Infrastructures 2020-2027 (object CERN), Bulgaria; Ministry of Education of China (MOEC) , Ministry of Science & Technology of China (MSTC) and National Natural Science Foundation of China (NSFC), China; Ministry of Science and Education and Croatian Science Foundation, Croatia; Centro de Aplicaciones Tecnológicas y Desarrollo Nuclear (CEADEN), Cubaenergía, Cuba; Ministry of Education, Youth and Sports of the Czech Republic, Czech Republic; The Danish Council for Independent Research | Natural Sciences, the VILLUM FONDEN and Danish National Research Foundation (DNRF), Denmark; Helsinki Institute of Physics (HIP), Finland; Commissariat à l’Energie Atomique (CEA) and Institut National de Physique Nucléaire et de Physique des Particules (IN2P3) and Centre National de la Recherche Scientifique (CNRS), France; Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR) and GSI Helmholtzzentrum für Schwerionenforschung GmbH, Germany; General Secretariat for Research and Technology, Ministry of Education, Research and Religions, Greece; National Research, Development and Innovation Office, Hungary; Department of Atomic Energy Government of India (DAE), Department of Science and Technology, Government of India (DST), University Grants Commission, Government of India (UGC) and Council of Scientific and Industrial Research (CSIR), India; National Research and Innovation Agency - BRIN, Indonesia; Istituto Nazionale di Fisica Nucleare (INFN), Italy; Japanese Ministry of Education, Culture, Sports, Science and Technology (MEXT) and Japan Society for the Promotion of Science (JSPS) KAKENHI, Japan; Consejo Nacional de Ciencia (CONACYT) y Tecnología, through Fondo de Cooperación Internacional en Ciencia y Tecnología (FONCICYT) and Dirección General de Asuntos del Personal Academico (DGAPA), Mexico; Nederlandse Organisatie voor Wetenschappelijk Onderzoek (NWO), Netherlands; The Research Council of Norway, Norway; Pontificia Universidad Católica del Perú, Peru; Ministry of Science and Higher Education, National Science Centre and WUT ID-UB, Poland; Korea Institute of Science and Technology Information and National Research Foundation of Korea (NRF), Republic of Korea; Ministry of Education and Scientific Research, Institute of Atomic Physics, Ministry of Research and Innovation and Institute of Atomic Physics and Universitatea Nationala de Stiinta si Tehnologie Politehnica Bucuresti, Romania; Ministerstvo skolstva, vyskumu, vyvoja a mladeze SR, Slovakia; National Research Foundation of South Africa, South Africa; Swedish Research Council (VR) and Knut & Alice Wallenberg Foundation (KAW), Sweden; European Organization for Nuclear Research, Switzerland; Suranaree University of Technology (SUT), National Science and Technology Development Agency (NSTDA) and National Science, Research and Innovation Fund (NSRF via PMU-B B05F650021), Thailand; Turkish Energy, Nuclear and Mineral Research Agency (TENMAK), Turkey; National Academy of Sciences of Ukraine, Ukraine; Science and Technology Facilities Council (STFC), United Kingdom; National Science Foundation of the United States of America (NSF) and United States Department of Energy, Office of Nuclear Physics (DOE NP), United States of America. In addition, individual groups or members have received support from: Czech Science Foundation (grant no. 23-07499S), Czech Republic; FORTE project, reg. no. CZ.02.01.01/00/22_008/0004632, Czech Republic, co-funded by the European Union, Czech Republic; European Research Council (grant no. 950692), European Union; Deutsche Forschungs Gemeinschaft (DFG, German Research Foundation) “Neutrinos and Dark Matter in Astro- and Particle Physics” (grant no. SFB 1258), Germany; ICSC - National Research Center for High Performance Computing, Big Data and Quantum Computing and FAIR - Future Artificial Intelligence Research, funded by the NextGenerationEU program (Italy).